Project

General

Profile

Actions

KaonLT/PionLT Meetings

KaonLT/PionLT collaboration analysis meetings. Please add your talk to the meeting ahead of time. Add the newest meeting to the top of the list!


2022 Meetings


Sept 28, 2022


Sept 21, 2022

  • Garth/Stephen notes mtg_22sep21.txt
  • Richard
    • Yield_6p2cl1.pdf Yield_10p6cl2.pdf Yield_10p6cl1.pdf
    • Lumi SHMS tracking is still showing the trend from 9/8. It was found that the lumi replay script is not updated with many of Ali and Richard's most recent changes. These changes are being applied and a set of new replays will be done.
    • Bill's cross-section code is broken into two major parts. The root analysis part, which takes all the yields, uncertainties, etc., is nearly complete with the connecting ltsep framework script finishing up. The second part, where many of the final plots and cross-sections are done, is a hard-coded mess. Richard is in the process of making this more dynamic. This aspect will take more time but the root analysis part is the bulk of the analysis anyways.
  • Ali
    • Ali_Sep21_Analysis.pdf
    • Heep Coin analysis for high Q2 KaonLT data (6.2, 8.2, 10.6 GeV beam energy) has now all the efficiencies/live times applied. The results for all 3 beam energies are looking very good. Data/SIMC ratios are agreeing very well (< 3%) and the shapes are matching up nicely. Next step is to finalize offsets for both low and high Q2 KaonLT data. Also proton absorption correction and target boiling needs to be applied.

Sept 8, 2022


Aug 31, 2022

  • Garth/Stephen Notes mtg_22aug31.txt
  • Richard
    • KaonLTMeeting8_31_22.pdf
    • Improvements to luminosity analysis. Things are getting within a few percent. There are still some obvious issues with dependency leaking into the cuts, especially for tracking which is being investigated further.
    • Uncertainties for the luminosities and heep are next up.

Aug 26, 2022


Aug 17, 2022

  • Garth/Stephen Notes mtg_22aug17.txt
  • Vijay
    • The summary of the Heepcheck study is attached here. Vijay_Aug17_2022.pdf
    • 3.6 and 3.9 GeV data do NOT have the angle offsets. This has been established in the study.
    • I've plotted the offsets vs SHMS central momentum. Plots are attached in the attached pdf file.
  • Richard
    • The 10.6 GeV heep analysis is going well but the broad momentum distributions are making offsets difficult at times.
    • I have decided to give a second pair of eyes to Ali's hcana vs simc calculation investigation. It is concluded that the broad peaks of the hcana are likely caused by the secondary particle calculation which is gathered from the spectrometer track info.
    • Corrections to the simc pmiss components have been added to the fortran code which corrects for a minus sign that is in hcana but missing from simc.
    • FullAnalysis_Offset_10p6.pdf
    • HcanavSimc_scatter_calc.pdf

Aug 4, 2022

  • Garth/Stephen Notes mtg_22aug04.txt
  • Vijay
    • The offset HeepCheck study at 3.9, 4.9 (KaonLT), 2.7, 3.6 and 4.5 (PionLT) GeV has been completed. Vijay_HeepCheckStudy.pdf
  • Richard
    • KaonLTMeeting_8_4_22.pdf
    • HeeP coin for 10p6 setting is looking pretty good. I still need to make some minor adjustments.
    • TLT for Lumi is looking much better. The yields also are looking better but cut adjustments need to be made.
    • HeeP singles have some odd issues that need to be resolved before offset analysis can begin.
    • Hodoscope calibrations are all done except 6p2.

July 14, 2022


June 28, 2022

  • Garth/Stephen Notes mtg_22jun28.txt
  • Richard
    • Aerogel XY plots vs NPE to check out the tray location and efficiencies.
    • Hodoscope and calorimeter scripts working. Hodoscope calibration looks good for 10p6 but needs improvement for other beam energies.
    • HGCer efficiency and plots added to efficiency script. Need to discuss with Vijay on specifics of cuts and efficiencies.
    • Started looking in detail at Bill's code. Some scripts and links are missing/broken.
  • Jacob
    • EDTM Study Report completed, with newer versions of the TLT equations EDTM_Study_Report.pdf
    • Graphical Guide for Event Types created to accompany the report, illustrating prescaling with EDTM Graphical_Guide_Evt_Type.pdf
    • Plan to update report file template with new equations and errors soon

June 14, 2022

  • Garth's Notes efficiencies_22jun13.txt
  • Richard
    • Fixed issues with simc script and 10p6 Heepcoin data. Issue with data was an offset of 40ns that was corrected halfway through runs.
    • Aerogel XY plots vs NPE to check out the tray location and efficiencies.
    • Efficiency plots for all production runs.
    • KaonLTMeeting6_14_22.pdf

May 31, 2022


May 17, 2022


May 3, 2022

  • Garth and Stephen's Notes mtg_22may03.odt efficiencies_22may03.odt
  • Ali
    • Fixed all the settings for high Q2 KaonLT data. Also slide on accessing spectrometer angles from CODA
    • KaonLT_May_3.pdf
  • Richard
    • Analysis of 10.6 GeV heep data. Need to scale data to see how they compare with SIMC for smaller offsets. KaonLTMeeting5_3_22.pdf
  • Vijay
    • The SIMC simulation study of the PinoLT data at Q2 = 0.38 and 0.42 GeV2 has completed (preliminary) for all settings.
    • An output pdf file of the 4.5 GeV, Q2 = 0.42 GeV2 and "center" setting is here Vijay_analysis_May3_2022.pdf .

April 19, 2022


April 06, 2022

  • Garth's Notes mtg_22apr06.txt
  • Ali
  • Richard
    • KaonLTMeeting4_6_22.pdf
    • The analysis of the 10.6 GeV heep coin/singles settings are still ongoing.
    • There were some issues getting the replays working for the singles runs and the analysis. Things finally were sorted out yesterday (and this morning).
    • The SHMS heep singles are presented. Overall, they look fine. There are still some things to sort out.
    • The HMS heep singles are still running.
  • Vijay
    • The simulation study of the 4.9 GeV KaonLT data has completed.
    • This energy has left, center and right settings data. All these settings look fine.
    • See the pdf for more info. Vijay_KaonLT_Meeting_Apri_06_2022.pdf

March 23, 2022

  • Garth's Notes mtg_22mar23.txt
  • Richard
    • 10.6 GeV heep coin was analyzed. Things look pretty good so far, still some work to be done. KaonLTMeeting3_23_22.pdf
    • Heep singles are next up. There have been issues replaying these runs but Ali may have fixed this issue with his parameter file changes.
  • Nathan
    • Did chack of the Mode 10 plots for HGC, was discrepancy between hardware and software caused by the variable phgcer_SampNSAT being 2 by default.
    • after changing back to 1 discrepancy resolved. Slides: Mode10Update_2022Mar23.pdf

March 9, 2022


February 23, 2022


January 26, 2022


January 12, 2022


January 5, 2022

  • Garth's Notes mtg_22jan05.txt
  • Stephen
    • Stephen was updating GitHub but UTIL_BATCH has some permissions issues
  • Richard
    • Update to KLT task schedule
    • Meeting between Richard and Vijay to discuss task breakdown Friday, January 7th at 11 am EST
    • EDTM study 2 will done sometime this weekend
  • Vijay
    • Breakdown of HEEP single analysis for low Q2 data
      • Shapes of SIMC and data are very off
      • It seems to be a systematic issue on Vijay's part because the same issue is being seen in PLT data
      • The main issue in my input file was the target thickness, which picked up by Dave in the group meeting. I corrected the target thickness and now it looks ok. See the attached pdf file below.
    • https://redmine.jlab.org/attachments/1376
    • https://redmine.jlab.org/attachments/1375
  • Nathan
  • Jacob

2021 Meetings

August 4, 2021

  • Richard
    • Update on the HEEP singles. The remainder of energy setting ratios vs SIMC were looked at over the range of angles. They look good (within 1% and < 0.5% uncertainty).
    • Issues with the COIN are seen but may be a replay issue.

July 21, 2021

  • Richard
    • Update on the HEEP singles. 10.6 GeV ratios vs SIMC were looked at over the range of angles. They look good (within 1%) but proper error bars are needed.

May 12, 2021

  • Richard
    • The currents were adjusted from 1-10 uA with little effect on the EDMTLT but proper changes to the CPULT. The replay bcm cut script developed by Sanghwa was also analyzed to see if it may be causing the issue but it had the same effects. There was also a small error in the code found where T.coin.pEDTM.tdcTimeRaw was incorrectly written as T.coin.pEDTM.tdcTime.

April 28, 2021

  • Richard
    • Lumi table was created for 2 sets of runs. Odd behavior seen ~30-45 uA with EDTM where is drops. Playing with the current cuts is the next step.

April 14, 2021

  • Richard
    • Updated Lumi carbon plots with uncertainties. PID and current cuts need to be played around with a bit to see if they improve things (as uncertainties don't explain any boiling-type behaviors).
    • Some PionLT carbon runs were looked at but there are still some issues (either cut or pathing related).

March 31, 2021

  • Richard
    • Luminosity studies
      • PionLT carbon runs
        • Still having issues getting decent results. I think I am just missing a definition somewhere that I need to track down
      • Uncertainties
        • Easy to do, I just need to put it into my script
    • Explicit PS calculations
      • I have gotten a rough script working for very simple runs (ie low rate, single spectrometer, calibration runs)
      • I am trying to expand to lumi singles runs, things get a bit more complex so I am just working out the trigger correlations
    • Pass 2 runs
      • DEF-files
        • I checked out the DEF-files for our combination of the nested files
        • There are a lot a files (~150!!!!) and may be hard to track so next week we should make a flowchart together
        • Working on my list of branches for the next replay
          • Mainly just working on my list of branches for the explicit PS calculations
          • Everyone should have a list of these branches by next mini-meeting (4/7/21)
  • Stephen
    • Updated DEF-File for the "full" replay
      • Can't remove NData brances, Mark says they won't take up too many resources anyway
      • Removed individual PMT info and raw info for many detectors, reduced file size by factor ~2
    • Updating DEF-files/scripts for calibrations as part of PionLT prep
    • Will focus on DEF-file/DEF-cut file updates over next week
      • Will investigate HeeP replays over rest of this week.

March 24, 2021


March 17, 2021

  • Richard
    • I have re-run the carbon lumi runs with the changes to the replay events (i.e. All_events_NO_EDTM). The results look pretty good (within a few percent). There are still some trends I need to investigate for certain runs and I need to do the analysis for the pionLT carbon lumi data.
  • Mini meeting March 24th @ 4:30 EDT

March 3, 2021

  • Richard
    • I found an issue with the replay scripts used where the EDTM was not being properly registered. There is a discrepancy between two replay scripts depending on what the Decode_master is set to (i.e. ALL_events vs ALL_events_NO_EDTM).
    • The explicit calculation of the prescale has hit a snag because of the differences between 6 GeV and 12 GeV. I contacted Mark Jones, who explained there are significant differences in the ADC gate between the two eras (which my own findings confirmed). I have contacted Brad for further input.

February 17, 2021


February 3, 2021

  • Richard
    • A table with the aerogel efficiency for kaons is being created. It will be expanded to other particles and detectors.

January 20, 2021

  • Richard
    • All KaonLT carbon runs have been analyzed. There are still some issues that seem to be related to the pre-scale values. I may need to investigate the EDTM and/or the COIN events further to correct for these.

January 6, 2021

  • Richard
    • Analyzed most of the luminosity scans. There are a few issues I am investigating still which seem to be run related.

2020 Meetings

December 2, 2020

  • Ali and Richard
    • Richard included the EDTM to Lumi analysis and found some old analysis of scalers by Mark Jones that he will compare. Ali and Richard met last week and will begin combining Ali's tracking studies with Richard's lumi analysis.

November 18, 2020

  • Carlos
    • He is investigating some of the EDTM rate dependence and sees a discrepancy between the total EDTM LT and CPU LT with a sudden dip at 20-30 kHz. More luminosity runs will need to be studied for a full conclusion.
  • Richard
    • Scaler analysis matches Carlos' code now. Overall significant improvements were made through some cut adjustments and incorrect variable definitions.
  • Stephen
    • Updated the hallc_replay_lt with parameter files for all coin peaks.

November 4, 2020

  • Richard
    • Scaler analysis is getting closer to an answer. An in-depth comparison of Carlos and Richard's analysis shows inconsistencies between the two.

October 28, 2020

  • Ali
    • Went over some of his tracking procedures for the report file implementation
  • Richard
    • Luminosity analysis continues. There are still some issues but things have improved a bit.
  • Stephen
    • Discussion of CT peak analysis, progress and outline of what's left to do with this
  • Vijay
    • Dry run for Mark's hallc meeting tomorrow (10/29)

October 14, 2020

  • Ali
    • Working through the REPORT_FILES template to get everything working and outputting proper information.
  • Richard
    • Updates on lumi scans. There are many issues from Scalers to Tracking that need to be resolved.
  • Stephen
    • He has included a demo for using the python analysis script in the README.

September 23, 2020

  • Richard
    • A quick lumi analysis after the Pass1 runs. The results were compared to previously analyzed runs (5154-5158). There was some change but nothing huge. Some other runs with the same energy settings were analyzed as well.
  • Stephen
    • Showed some slides from u-channel meeting of omega and proton MM plots. Then discussed some ideas for dealing with the hole in the HGCer.
  • Petr
    • Data fitting algorithm based on E.H. Bellamy paper. Extracts PMT spectra from ROOT file and makes a 1D projection histogram.

September 9, 2020

  • Richard
    • Still working on run list, GUI checks
  • Stephen
    • New slim replay scripts and full replay scripts updated
  • Ali
    • Him and Stephen found an issue (already aware of) with zeros in the parameter files using Mark's hcana changes
    • Need to be vigilant of crashed replays because this is probably the issue
  • Vijay
    • Nearly finished with HGcer calibrations

August 26, 2020

Links to talks-> Ali Vijay

  • Richard
    • New tasks were created based on the previous discussions on objectives going forward. In the process of checking run sheets against the run list and camera/EPICS.
  • Stephen
    • Updated on RF times. There were a few issues with some runs, but overall most are finished.
  • Ali
    • Hodoscope calibrations look similar to before. There is still the issue of the double bands in the time walk. This is ok because of the wide windows we still have. (See hodoscope calibration section)
  • Vijay
    • Updated the SHMS HGCher analysis. The calibrations are coming along, but a little more work needs to be finished up.

July 29, 2020

Links to talks-> Richard Stephen

  • Stephen
    • Updates on scripts to set pedestal defaults and AdcTdc offsets after hcana change
  • Richard
    • Updated plan and schedule for the analysis going forward

July 8, 2020

Links to talks-> Vijay Ali


June 24, 2020

Links to talks-> Vijay Richard


June 9, 2020

Links to talks-> Vijay Ali Richard


May 27, 2020

Links to talks-> Vijay Ali


May 13, 2020

Links to talks-> Stephen - Python analysis scripts Richard Ali Vijay


April 29, 2020

Links to talks-> Richard Stephen - Python code planning Mark's hcana tracking update


April 15, 2020

Links to talks-> Richard


April 1, 2020

Links to talks-> Stephen - RF timing Richard


March 18, 2020

Links to talks-> Richard's EIC Temple Draft 2


March 11, 2020

Links to talks-> Ali Richard's EIC Temple Draft 1


Feb 26, 2020

  • Peter's changes to hcana should be coming soon. Richard will check with Mark when he is down there for an update.
  • Need to incorporate the latest BCM updates from Dave Mack.
  • Richard should look at the luminosity scalar analysis before and after the BCM changes once he gets his script working again.
    • He is checking with Mark about this while he is down in jlab.

Feb 10, 2020


Jan 22, 2020

Links to talks-> Richard's Hall C Winter Talk Draft 1


Jan 15,2020

Links to talks-> Richard's Hall C Winter Talk Outline


Jan 8,2020

Links to talks-> Richard

  • Richard
    • PID efficiency script is coming together
      • Tighter cuts and cuts on timing, beta, etc. will be needed for cleaner electron cuts
      • Hcal cut should be changed to cut around the mean normalized value instead of cutting above or below it only. This will assure electrons aren't lost.

2019 Meetings


Dec 17,2019

Reminder: Mark Jones' commissioning analysis meetings are bi-weekly and announced by Mark

  • Stephen
    • Number of passes finished for HMS calorimeter
      • Still wiggles in all passes
    • Finished his report on calibrations, DC and Calorimeters
    • Put together a list of people's jobs from Regina over the next couple of months, Regina Analysis Tasks
      • CUA will be putting together a similar list
  • Richard
    • Fixed one error issue with SHMS calorimeter calibration code (discussed Dec 10th meeting), but still have one error
    • PID efficiency script plots are complete, but still some questions he needs to work out before a final version
      • Eventually need to include tracking into PID efficiencies
    • Pete brought up the question for a number of multi-hits we are getting
      • (Dead Time)+(Trigger efficiency)-EDTM

Dec 10,2019

Links to talks-> Richard

  • Garth and Stephen were in the most recent SIDIS meeting
    • Wit had lots of hcana changes but they seem to be critical to our work
      • Changes such as new detector classes
    • Hopefully Wit will present these in the next commissioning meeting
  • Stephen improved wiggles but still prevalent
  • Richard finished HGCer calibrations, he concluded that the online parameters are sufficient
    • He also has been looking at SHMS calorimeter calibrations but has had issues running the scripts. Most likely a pathing issue.
    • The tasks page has been updated with a section for efficiencies.
      • A PID efficiency script is currently being developed and hopefully should have some plots for our next meeting.

Nov 26,2019

Links to talks-> Richard

  • Stephen
    • DC SHMS, small-angle has poor performance
      • Tighter cuts or PID (currently only looks for electrons)
    • Vardan said calorimeter wiggles are an ongoing issue with no current solution
  • Richard
    • The calibrations are nearing completion with only the HG Cherenkov and SHMS calorimeter yet to be finished. There are also smaller issues with some DC and HMS calorimeter runs.
      • Richard will pick up looking at the HGCer calibrations from Vijay's scripts.
      • Richard will also be looking at the SHMS calorimeter and with the suggestion of getting some negatively polarized SHMS runs from Hem.
    • The HMS Cherenkov is finished with the conclusion that the calibrations from the online runs are good
    • We can begin looking forward towards efficiencies.
      • Garth suggested that looking at PID may be an advantageous first steps. Precise definitions of these PIDs will need to be logged.
      • Clear notes on efficiencies steps will be needed. They will need to be redefined multiple times in multiple ways.
  • Hem
    • Has been looking at the calorimeter as well.
    • He has seen the second band in the spectra for some runs, we suggested further looks into HMS hodoscope and subsequently reference times. A recalibration will need to be performed.

Nov 19,2019

Links to talks-> Stephen Richard

  • Richard
    • Finished up TDC cut paper going through the procedure (found here)
    • Needed to add Carlos pre-trigger designs to an appendix for the paper which clarifies the introduction section
  • Stephen
    • Issues with DC calibration for only a few runs during each run period
    • "wiggles" in HMS calorimeter calibrations for some runs
      • few runs where HMS cal totally failed (which was dependent on initial parameters)
      • Spring looked worse than the rest (surprisingly), but this may be because online calibration wasn't great

Nov 5,2019

Links to talks-> Stephen

  • Timing update (Stephen)
    • Reference times between Richard and Stephen are very similar and the GitHub is updated accordingly
    • Richard is going to add the trigger time cuts once the fall runs are finished (by end of week)
    • First pass of detector reference time parameter files finished
  • Peter's hcana code (Richard and Stephen)
    • Peter's code does not remove ghost bands (already known), but now Stephen is seeing a strang horizontal band for just SHMS hodoscope 2D plots
      • Richard has emailed Peter about the issue with no response yet (follow-up soon)
    • This means we still need the reference time amplitude cut with Peter's code but using just the amplitude cut with Mark's latest version of hcana should be sufficient for calibrations.
  • Calibrations were decided to be done once per shift
  • Next meeting schedule...
    • Small meetings Nov 12 and Nov 19 (at 3pm EST) so Richard and Stephen can iron out the timing and calibrations
    • November 26 (at 3pm EST) will be a larger meeting.
      • This is the rough deadline we made for finishing up calibrations.
      • This should be sufficient for all detectors (although hodoscope and HG which may persist). The ultimate goal is for all calibrations to be finished up before Winter break!!

Oct 22,2019

Links to talks-> Stephen Richard

  • SHMS Reference Time Amplitude Cuts(Stephen)
    • Extra peaks in HG from cross-talk and "ghost-bands" due to ADC screw up
      • cuts applied on SHMS ADC ref time amplitude (greater than zero) done before any further event selection and histogram filling
    • This is a temporary fix to "ghost-bands" and cross talk
      • Although for HG there is still a small peak. The hodoscope is fine.
    • The time window is easier to set now.
    • DC has been fine (since it uses TDC)
      • DC calibrations should be good to complete, other detectors need more fixes (i.e. the remaining "ghost-band")
    • Hodoscope still bad but this might be because of that "ghost-band", needs to be further investigated
  • Peter's hcana fixes for reference time (Richard)
    • Peter implemented changes to hcana
    • Two major issues...
      1. Amplitude goes to null and integral is low or negative which causes a rate-dependent efficiency.
        • Peter has used the average pulse value acquired over a large sample of runs for each detector.
      2. Timing offset between EL-REAL and 3/4 is 130 ns which causes junk tracks at high SHMS rates (~6% of the time)
        • Peter created a routine to find the best offset by means of matching TDC and ADC
        • Richard has been checking for bugs in the code.
          • It seems to be ready to implement, Richard emailed Peter for confirmation on the implementation in our group directory
  • Calibration of HMS detectors when the timing is agreed upon by Stephen and Richard.
    • Besides hodoscope which still has an issue with the script
    • Stephen and Richard are having a meeting late next week to determine time window values
      • These will be relatively final, but on the wider side if adjustments need to be made in the future

Sep 24,2019

Links to talks-> Stephen Richard


Sep 10,2019

Links to talks-> Stephen Ali Richard


August 26, 2019

Links to talks-> Stephen Ali Vijay Richard


June 21, 2019

Links to talks-> Hallc Talk Outline


June 14, 2019

Links to talks-> Hallc Talk Outline Mireille Update-Mireille Vijay


June 4, 2019

Links to talks-> Stephen Nathan Vijay Richard Hallc Talk Outline Mireille


May 21, 2019

Links to talks-> Richard Mireille? Stephen Nathan Vijay


May 14, 2019

Links to talks-> Richard Carlos' slides Mireille Stephen


April 30, 2019

Links to talks-> Richard Richard's slides in wiki Brad Computing Stephen


February 4, 2019

Links to talks-> Richard Stephen


2018 Meetings


August 21, 2018

Ongoing Tasks-> #381 #382 #383 #404 #405 #406


August 8, 2018

Ongoing Tasks-> #381 #382 #383 #404 #405 #406


August 3, 2018

Ongoing Tasks-> #381 #382 #383
Links to talks->


July 11, 2018

Ongoing Tasks-> #381 #382 #383 #384

Action items:

  • Update run plan, e.g. add prescales (Garth)
  • Check on details of: magnet cycling procedure, minimum angles, MCC fringe field steering, beam energy measurements, optics checks (Tanja)
  • Identify additional run coordinators (Tanja)
  • Encourage collaborators to take shifts, e.g. FIU (Pete)
  • Draft email about shifts etc. to collaboration (Tanja)
  • Develop diagnostic scripts and kaonLT specific (Richard, Ryan)
  • Continue documentation development, e.g. checklist, shift taker instructions, HowTo or link to generic Hall C HowTo, task list, web page (Richard, Garth)

Discussions:

  • SHMS Aerogel Cherenkov: tray has been changed to SP-11 on 10 July, nominal time for change is 8 hours, one tray can stay in the hall to minimize overhead
  • HGC: waiting for re-installation
  • Trigger checkout: main items to keep in mind: fADC thresholds, timing HMS/SHMS - leading arm, possible rate limitations
  • Shifts
  • Diagnostics scripts and documentation
  • Personnel schedules

June 28, 2018

Links to talks-> Richard

Action Items:

  • Develop and post shift schedule. Send out request. Find out who is KAONLT liaison (Tanja)
  • Contact potential run coordinators. (Tanja)
  • Aerogel tray exchange for start of experiment and coordination of future exchanges (Vladimir)
  • Update runplan and prepare documents needed for shift crew, e.g. run sheets (Garth)
  • Set up a web page to collect and link from Redmine - How-To documents, initial calibrations, etc. (Garth, Richard)
  • Adapt luminosity scan script to calculate kaon yields by run number for use during experiment by shift crew - has to be simple to use. (Richard)
  • Develop trigger/detector checkout plan (Tanja, initial draft)
    • Are all experts available at run start, experts needed: coincidence timing , Harp Scans, BCM calibrations, ...
  • Find out from JLab what the anticipated settings are for limit switches for both spectrometers and for separation angle in between them. Also find out what magnet cycling procedure should be used for L/T quality separation data (Tanja)
  • Find out if all aerogel trays needed for experiment can be stored inside Hall C to avoid having to open truck ramp (Vladimir)

Analysis Updates:

  • Ryan
    • resolving differences between his analysis and Richard's
  • Richard
    • Discussed with Mark Jones and is cross checking his January luminosity scan analysis against Mark's. Realized that BCM calibrations (gains, offsets etc) are different - currently working on fixing that, presents current results - low curent point is ~10% below unity, discussion that this seems to indicate that BCM offsets are the issue as low current points would be most sensitive, is planning to also double-check the BCM calibration procedure - had discussion with Dave Mack at JLab
    • Presents a script developed for luminosity scans that determines yields for a run number input on command line. Discussion this could be good to adapt for kaon counting during experiment.

May 31, 2018

Links to talks-> Samip


May 17, 2018

Links to talks-> Richard, Ryan, Samip

  • Luminosity scans (Richard, Ryan)
    • Carbon scan from April seems consistent with no boiling, LH3 and LD2 show deviation from unity, should check effect of EDTM events on efficiencies
    • Carbon scans from January (low rates) have large deviations from unity, in particular behavior at higher currents (>40uA), e.g. cpu dead time, that is not yet understood
    • discussion about pre-scale factors
  • Discussion about Richard's script extracting and printing information from report files and trees including histograms
    • Add option for looking at yields with cuts on cointime, target window, and missing mass
    • Add option to look at real peak vs. background level
  • H(e,e'p) kinematics investigation for fall run (Samip)
    • Optimal range is between electron angles of 17-20 degrees, time required ranges from ~1 hour at 17 degrees to ~3 hour at 20 degrees, other angles may be of interest for disentangling charge and current dependent effects
  • Action Items
    • Luminosity scans at low rates from January (Richard)
    • Luminosity scans from April with corrected efficiencies, e.d. DT (Ryan)
    • H(e,e'p) and H(e,e') kinematics for upcoming Kaon LT run (Samip)
    • Kaon LT detailed run plan (Garth, Pete, Tanja)
    • Preparations, e.g. report files, diagnostic histograms and real time monitoring, shift worker instructions, optimization of trees (all)

May 10, 2018

Links to talks-> Hem, Richard, Ryan, Samip

  • Luminosity Scans January 2018 (Richard)
    • Tracking-based analysis for Runs 1415-1423, .. for now focus on HMS, typical HMS S1X rates are <
    • Cuts applied: PID for electrons (Cer, Cal), acceptance (delta)
    • Efficiency calculation was investigated in detail and applied using scalers from trees
    • Results for carbon (normalized to 30uA) show large variation - need to understand
    • Next steps:
      • Double check calculation of computer lifetime, look for differences in TRIG and PRETRIG scalers
      • Check stability of pre-scale (Ryan saw changes within runs)
      • Check if detector calibrations for runs are implemented correctly
      • Analyze with Ryan's cuts
  • Luminosity Scans April 2018 (Ryan)
    • Confirms Richard's earlier finding that scaler-based analysis is not reliable and that it's best to analyze event-by-event
    • Tracking-based analysis for runs 2073-2095
    • Cuts applied: PID for electrons (TOF, Cal, Chi2), acceptance (delta and angles), BCM4a
    • Included most recent detector calibrations in replay of these runs
    • Efficiencies from report file mainly, may need to update tracking efficiency implementation
    • Found that calculations of computer deadtime requires different set of scalers for runs from January to those from April 2018
    • Results show carbon yields consistent with unity within error (0.5% error bars), and boiling for hydrogen and deuterium
    • Next steps:
      • Check tracking efficiency
  • Global comments on Luminosity Scans (so far)
    • April 2018 scans show hydrogen and deuterium target boiling on level of 3-4%, Carbon is consistent with no boiling, as expected - based on description, new targets are expected to have less boiling than those used previously, e.g. in FPI2. Double check this.
    • Efficiencies in report files need to be updated, e.g. CPU deadtime (Richard)
    • Check if BCM4a is the one closest to the target - get a map of the hall (Richard, Ryan)
    • Uncertainties on current
    • Detector calibrations
    • HMS vs. SHMS
  • eK+ coincidence time from SIDIS Spring 2018 run (Samip)
    • Worked with Mark Jones on coincidence time calculation - now available in hcana (see Samip's 10 May 2018 email for more details)
    • runs used in this analysis: 3423 with Q2=3.1 GeV2 and W=2.78 GeV
    • Coincidence time spectrum shows peak separations >2 ns corresponding to different accelerator frequency
    • Missing mass spectra: Kaon peak clearly visible, also see pions even after cuts
    • Found that pions appear localized in phase space - seems to be due to end point in kinematic phase space
    • Next steps:
      • Look for a run with kaon momentum closer to 3 GeV/c (aerogel threshold is ~3 GeV/c for kaons for n=1.015)
  • Yield normalized ratios from SIDIS (Hem)
    • Yield ratios of data to MC match well in shape, but normalization seems not quite right
    • Next steps:
      • Check if efficiencies, currents, run time etc. are taken into account correctly

April 26, 2018

Links to talks-> Richard, Ryan, Samip

  • Discussion luminosity scan analysis (earlier scan) - Richard
    • Shown is a scaler-based analysis using scalers from report file and from ROOT tree - the values are different, why that is has to be checked
    • Notes on efficiencies in report file:
      • DT (EDTM) is not electronic deadtime. Instead use for now 50ns* trigger rate for HMS and 100ns* trigger rate for SHMS
      • Cherenkov efficiency seems low (check if that is due to PID, thresholds, pi/e, etc)
      • Tracking efficiency needs to be checked
      • 3/4 trigger efficiency seems ok
  • Discussion luminosity scan analysis (later scan) - Ryan
    • Fewer carbon runs than earlier scan
    • Shown is a tracking-based analysis using data from the ROOT trees
    • First results for carbon, LH3, and LD2, but efficiencies need to be checked
  • SIDIS analysis - Hem
    • Shown is a data/MC comparison for different kinematic settings. Agreement in shape is very good. Normalization has yet to be applied.
    • Discussion about agreement in z and apparent offset - relation of this offset with spectrometer offset
  • Analysis of exclusive events in SIDIS - Samip
    • Shown is a look at the first kaon events in exclusive limit of SIDIS data
    • Ongoing work on implementing coincidence time into tree
  • Discussion of run plan, and in particular possible low Q2/low energy
    • Assume that kaon low Q2=0.4 already included - have kinematics with specified beam energies, if needed
    • For pion form factor and LT, have two possible settings that can use the prescribed beam energies: Q2=1.45 and 0.35 GeV2 - prioritize towards the higher Q2 as not as sensitive to offsets
    • Garth will put this together in a draft email

March 29, 2018

Links to talks-> Richard, Samip

  • Details of adding points to fall run plan - will be ready sooner, likely in week of 4/1 (Garth, Samip, Tanja)
  • Luminosity scans (Ryan, Andres, Richard, Pete)
  • Macros, SIDIS analysis (Hem, Anashe)
  • Elastic data (TBD)

February 22, 2018

Links to talks-> Andres

  • Discussion of Andres' slides on elastic W calculation in hcana - initial issue with negative W values seems to be due to outdated information in standard.kinematics file. Additional investigation ongoing.
  • Brief description of Samip's work - working on coincidence timing decoding with JLab staff and re-optimization of kinematics
  • Discussion of Richard's slides on run plan kinematics and preparations -
    • everyone agrees that in general want to have as many L/T cross section separations as possible in a robust run plan balancing statistics vs. systematics
    • before actual run may want to do a 3D optimization in all kinematic variables to get the best possible delta(epsilon) for the L/T separations
    • for each setting will need 1-2 shifts for calibrations - need to negotiate with other experiments in run group
    • kinematics plots W-Q2 show that still need to do some optimization
    • run sheets: include raster size, update coincidence count labeling, and carbon target labeling
  • Discussion of next meeting. Garth at JLab 8-16 March - might have meeting during that time.
  • Action Items:
    • Post slides/notes from this meeting and procedure to set up hcana in Redmine (Tanja)
    • Update run sheet and kinematic optimization for W/Q2 (Garth)
    • Circulate elastic calculations for kinematic settings (Samip)
    • Continue investigation of elastic data and run plan kinematics (Andres/Richard)
    • Test software installation instructions (Pete)
    • Draft email to JLab management about need for small angle beam pipe - need to have HMS minimum angle of less than 13 degrees for kaon L/T separations (Tanja)

Updated by Jacob Murphy 3 days ago · 396 revisions